OpenStudy (anonymous):

how do i put 0.01666666666 into a fraction

OpenStudy (freckles):

Let x=0.01666666666666..... Multiply both sides by 100 100x=1.666666666666666666............. Now do 100x-x=1.6666666666666......-0.016666666666666.... or just 100x-x=1.66-0.01 since the other 6's cancel

OpenStudy (freckles):

And end the you can might still have a decimal but you can multiply by multiplies of 10 on bottom and top to get rid of it ...

OpenStudy (aum):

5 / 3 = 1.6666.... 0.5 / 3 = 0.166666.... 0.05 / 3 = 0.016666..... 0.05 / 3 = 5 / 300 = 1 / 60

OpenStudy (freckles):

I will finish my way since you already have the answer... \[\text{ Let } x =0.016666666666666... \\ \text{ So } 100x=1.66666666666666666...\\ 100x-1x=1.666666666...-0.01666666... \\ 99x=1.66-0.01 \\99x=1.65 \\ x=\frac{1.65}{99} \\ x=\frac{1.65}{99} \cdot \frac{100}{100} \\ x=\frac{165}{9900}\\ \text{ This fraction can be reduced } \\ x=\frac{165}{165 \cdot 60}=\frac{1}{60}\]